OpenStudy (anonymous):

the bse of a triangle is square root of 18 cm and its height is square root of 8 cm. find its area

OpenStudy (anonymous):

Area of triangle = 1/2 x base x height \[1/2\times \sqrt{18}\times \sqrt{8}\] Area = 6 cm squared

OpenStudy (anonymous):

Don't forget to fan :P